OpenStudy (bloomlocke367):

Suppose Q and R are independent events. Find P(Q and R). P(Q)=0.41 P(R)=0.44

OpenStudy (bloomlocke367):

@rafrasan

OpenStudy (anonymous):

I do not understand it and I have to go soon, sorry :(

OpenStudy (bloomlocke367):

Okay...

OpenStudy (bloomlocke367):

@iambatman

OpenStudy (bloomlocke367):

@myininaya

OpenStudy (bloomlocke367):

@saynabahmed13

OpenStudy (anonymous):

I can help

OpenStudy (bloomlocke367):

Okay, good :) what exactly is it asking me to do?

OpenStudy (anonymous):

Find P(Q and R)

OpenStudy (anonymous):

multiply the probabilities, (.41)(.44)

OpenStudy (bloomlocke367):

so I just multiply the two?

OpenStudy (bloomlocke367):

0.1804?

OpenStudy (anonymous):

yup

OpenStudy (bloomlocke367):

why?

OpenStudy (anonymous):

thats are answer

OpenStudy (anonymous):

because you multiply P(Q)=0.41 P(R)=0.44

OpenStudy (bloomlocke367):

but why do you multiply them?

OpenStudy (bloomlocke367):

@saynabahmed13

OpenStudy (anonymous):

(1) P (Q and R) = P(Q) x P(R) = 0.41 x 0.44 = 0.1804 [answer b] (2) P (S or T) = P(S) + P(T) = 8/9 + 1/10 = 80/90 + 9/90 = 89/90 [answer a] (3) P (S or T) = P(S) + P(T) = 6/11 + 1/10 = 60/110 + 11/110 = 71/110 [answer c]

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (anonymous):

there is and explanation

OpenStudy (bloomlocke367):

ohhh, okay. can you help with another?